# hyperExp head to column names, separator = comma, centered tabular = —c—c—c—c—c—c—c—c—c—c—, table head =     FW   FW EleSym   AGM   AGM EleSym Error Rel. time S(%) Rel. time S(%) Rel. time S(%) Rel. time S(%) , late after line =

### v3

@@ 5.1.1 A comparison between Renegar’s AGM, Algorithm 2 and DDS
In this series of experiments we proceed as follows. We fix the values of $n$ and $k$ in (5.1) and then we generate $30$ normally distributed points in $\mathbb{R}^{n}$. These are the $c$’s we would like to project onto $\mathbb{R}_{+}^{n,(k)}$. For each generated point $c$ we check if the minimum eigenvalue of $c$ with respect to $\mathbb{R}_{+}^{n,(k)}$ and $e=(1,\ldots,1)$ is greater than $-10^{-4}$. If this happens, then $c$ is deemed to be too close to the cone, so we discard it and generate a new point.
@@
We consider the objective function value obtained by DDS as the baseline to which we compare the performance of the other algorithms. The results are described in Tables LABEL:table:hyper_high_deg and LABEL:table:hyper_low_deg. We now explain the meaning of the data. For example, consider the first line in Table LABEL:table:hyper_hig_deg_10_1, so that the “Error” column indicates “$10\%$”. Roughly speaking, for this line, we checked how much time does each one of the 4 tested methods need to get a solution that has a value that is within $10\%$ of the function valued obtained by DDS.
@@
$f_{\text{FW}}^{i,j}\leq f_{\text{DDS}}^{i}\times\left(\frac{100+E}{100}\right),$ (5.6)
holds and the minimal eigenvalue of corresponding primal iterate (the $x_{k}$ in Algorithm 2) is at least $-10^{-8}$ (i.e., $x_{k}$ is sufficiently close to being feasible). Then, we record the ratio $\frac{t_{\text{FW}}^{i,j}}{t_{\text{DDS}}^{i}}$ and in the column “Rel. time” we register the average of these ratios together with their standard deviation. This average is what we henceforth call the mean relative time. In the “S” (for Success) column, we indicate the percentage of instances for which the method was able to find a solution that satisfies (5.6) within the feasibility requirements. The mean relative time is computed using successful instances only.
@@
For example, for the first line in Table LABEL:table:hyper_hig_deg_10_1 (which corresponds to $n=10$, $k=1$), the entry “$0.25\pm 0.08$” under “FW” means that, on average, our proposed method was able to find a solution whose objective value is within $10\%$ of the value found by DDS using $0.25\%$ of the time that DDS needed to find $f_{\text{DDS}}^{i}$. The corresponding standard deviation was $0.08$. It also succeeds for all $30$ instances, i.e., for each one of the instances there was at least one iteration that satisfied (5.6) with $E=10$. The data in the other columns “FW EleSym”, “AGM” and “AGM EleSym” have analogous meaning.
@@
Finally, for the experiments described in Tables LABEL:table:hyper_high_deg and LABEL:table:hyper_low_deg we provide extra statistics about the FW and FW EleSym approaches, see Tables LABEL:table:hyper_high_deg_det and LABEL:table:hyper_low_deg_det respectively. In those tables, we provide the average number of iterations and the absolute time required to reach each error target. Also, in the column “$\left\|x_{k}-x_{\text{DDS}}\right\|_{\infty}$” we provide the average of the distances (in the infinity norm) to the solutions obtained by the DDS solver. For example, in Table LABEL:table:hyper_hig_deg_10_1_det, in the block associated to FW, we have “$7.7\pm 2.9$” in the first row of the “Iterations” column. This indicates that considering all the $30$ points generated in Table LABEL:table:hyper_hig_deg_10_1, FW required an average of $7.7$ iterations to reach a solution with $E=10\%$ and the standard deviation was $2.9$. The average absolute time in seconds to obtain such a solution was $2.49\times 10^{-3}$ and the standard deviation was $8.14\times 10^{-4}$. Here we use “absolute time” to distinguish from the relative time described previously. Measuring the distance to the solution obtained by the DDS solver we obtained an average of $2.58\times 10^{-2}$ and standard deviation of $1.92\times 10^{-2}$.

### v4

@@ 5.1.1 A comparison between Renegar’s AGM, Algorithm 2 and DDS
In this series of experiments we proceed as follows. We fix the values of $n$ and $k$ in (5.1) and then we generate $30$ normally distributed points in $\mathbb{R}^{n}$. These are the $c$’s we would like to project onto $\mathbb{R}_{+}^{n,(k)}$. For each generated point $c$ we check if the minimum eigenvalue of $c$ with respect to $\mathbb{R}_{+}^{n,(k)}$ and $e=(1,\ldots,1)$ is greater than $10^{-4}$. If this happens, then $c$ is deemed to be too close to the cone, so we discard it and generate a new point.
@@
We consider the objective function value obtained by DDS as the baseline to which we compare the performance of the other algorithms. The results are described in Tables LABEL:table:hyper_high_deg and LABEL:table:hyper_low_deg. We now explain the meaning of the data. For example, consider the first line in Table LABEL:table:hyper_hig_deg_10_1, so that the “Error” column indicates $10\%$”. Roughly speaking, for this line, we checked how much time does each one of the 4 tested methods need to get a solution that has a value that is within $10\%$ of the function valued obtained by DDS.
@@
$f_{\text{FW}}^{i,j}\leq f_{\text{DDS}}^{i}\times\left(\frac{100+E}{100}\right)$ (5.6)
holds and the minimal eigenvalue of corresponding primal iterate (the $x_{k}$ in Algorithm 2) is at least $10^{-8}$ (i.e., $x_{k}$ is sufficiently close to being feasible). Then, we record the ratio $\frac{t_{\text{FW}}^{i,j}}{t_{\text{DDS}}^{i}}$ and in the column “Rel. time” we register the average of these ratios together with their standard deviation. This average is what we henceforth call the mean relative time. In the “S” (for Success) column, we indicate the percentage of instances for which the method was able to find a solution that satisfies (5.6) within the feasibility requirements. The mean relative time is computed using successful instances only.
@@
For example, for the first line in Table LABEL:table:hyper_hig_deg_10_1 (which corresponds to $n=10$, $k=1$), the entry $0.25\pm 0.08$” under “FW” means that, on average, our proposed method was able to find a solution whose objective value is within $10\%$ of the value found by DDS using $0.25\%$ of the time that DDS needed to find $f_{\text{DDS}}^{i}$. The corresponding standard deviation was $0.08$. It also succeeds for all $30$ instances, i.e., for each one of the instances there was at least one iteration that satisfied (5.6) with $E=10$. The data in the other columns “FW EleSym”, “AGM” and “AGM EleSym” have analogous meaning.
@@
Finally, for the experiments described in Tables LABEL:table:hyper_high_deg and LABEL:table:hyper_low_deg we provide extra statistics about the FW and FW EleSym approaches, see Tables LABEL:table:hyper_high_deg_det and LABEL:table:hyper_low_deg_det respectively. In those tables, we provide the average number of iterations and the absolute time required to reach each error target. Also, in the column $\left\|x_{k}-x_{\text{DDS}}\right\|_{\infty}$” we provide the average of the distances (in the infinity norm) to the solutions obtained by the DDS solver. For example, in Table LABEL:table:hyper_high_deg_detLABEL:table:hyper_hig_deg_10_1_detLABEL:table:hyper_hig_deg_10_1_det, in the block associated to FW, we have $7.7\pm 2.9$” in the first row of the “Iterations” column. This indicates that considering all the $30$ points generated in Table LABEL:table:hyper_hig_deg_10_1, FW required an average of $7.7$ iterations to reach a solution with $E=10\%$ and the standard deviation was $2.9$. The average absolute time in seconds to obtain such a solution was $2.49\times 10^{-3}$ and the standard deviation was $8.14\times 10^{-4}$. Here we use “absolute time” to distinguish from the relative time described previously. Measuring the distance to the solution obtained by the DDS solver we obtained an average of $2.58\times 10^{-2}$ and standard deviation of $1.92\times 10^{-2}$.